# Jiatong Li [^1] footnotemark: 1    Junxian Li3    Weida Wang2,4    Yunqing Liu1    Changmeng Zheng1 Dongzhan Zhou [^2] footnotemark: 2    Xiao-yong Wei1    Qing Li1

### v3

@@ Abstract
Recently, Large Language Models (LLMs) have shown great potential in natural language-driven molecule discovery. However, existing datasets and benchmarks for molecule-text alignment are predominantly built on a one-to-one mapping, measuring LLMs’ ability to retrieve a single, pre-defined answer, rather than their creative potential to generate diverse, yet equally valid, molecular candidates. To address this critical gap, we propose Speak-to-Structure (S2-Bench), the first benchmark to evaluate LLMs in open-domain natural language-driven molecule generation. S2-Bench is specifically designed for one-to-many relationships, challenging LLMs to demonstrate genuine molecular understanding and generation capabilities. Our benchmark includes three key tasks: molecule editing (MolEdit), molecule optimization (MolOpt), and customized molecule generation (MolCustom), each probing a different aspect of molecule discovery. We also introduce OpenMolIns, a large-scale instruction tuning dataset that enables Llama-3.1-8B to surpass the most powerful LLMs like GPT-4o and Claude-3.5 on S2-Bench. Our comprehensive evaluation of 28 LLMs shifts the focus from simple pattern recall to realistic molecular design, paving the way for more capable LLMs in natural language-driven molecule discovery.
1 Introduction
@@
Molecule discovery plays a pivotal role in various scientific research fields, from pharmaceuticals (Keiser et al., 2010) to materials science (Higuchi et al., 2023). Traditionally, this is a trial-and-error process (Ekins, 2024) that requires extensive experimentation and data analysis (Mattern & Grosser, 2023), often taking over a decade to bring a new drug candidate to market (Lee et al., 2018).
@@
Recently, Large Language Models (LLMs) have demonstrated great potential in molecule discovery (Edwards et al., 2021; 2022) by leveraging their powerful language understanding and robust reasoning abilities (OpenAI, 2023). Molecules can be represented as SMILES strings (Weininger, 1988), which allows LLMs to process molecules as textual strings, effectively bridging the gap between molecular structures and natural language. Meanwhile, by aligning molecules with textual descriptions (Edwards et al., 2022), LLMs can assist chemists in interpreting chemical knowledge, suggesting structural modifications, and predicting compound properties (Li et al., 2024a; Zhang et al., 2024), thereby significantly streamlining the molecule discovery process.
@@
While the integration of LLMs holds immense promise, a significant challenge lies in the limitations of current datasets and benchmarks. For example, existing datasets for aligning molecules and texts like ChEBI-20 (Edwards et al., 2022) and PubChem324K (Liu et al., 2023), are constructed based on a one-to-one mapping assumption, where each textual description is linked to a single, predefined target molecule due to the convenience of acquiring ground truth data from existing databases. However, the one-to-one mapping presents a key limitation: it does not fully align with the nature of real-world molecule discovery. In practice, multiple distinct molecules can often share the same desired properties or biological activity Petrone et al. (2012). For instance, a particular pharmaceutical effect is rarely unique to a single compound, and a material with a specific physical property, like tensile strength, can be realized through various molecular structures. This mismatch between the evaluation paradigm and the reality of molecule discovery casts doubt on whether these datasets and benchmarks genuinely assess an LLM’s capacity for creative molecular design, or if they instead inadvertently encourage models to rely on memorization and pattern-matching.
@@
To bridge this critical gap, we introduce a novel benchmark, Speak-to-Structure (S2-Bench), the first benchmark designed to evaluate the open-domain natural language-driven molecule generation capabilities of LLMs. S2-Bench is meticulously structured into three primary tasks, as shown in Figure 1, each targeting a specific, real-world capability essential for drug and materials science:
@@
Molecule Editing (MolEdit) focuses on testing an LLM’s ability to perform precise, localized structural modifications while preserving main structures. In drug discovery, this is analogous to Lead Optimization Barcelos et al. (2022), which edits the lead compound to create new variants. This requires a deep understanding of chemical grammar, including valency, stereochemistry, and ring systems, ensuring that edits could create valid molecules.
@@
Figure 1: Task illustration of S2-Bench for open domain natural language-driven molecule generation. In contrast to text-based target molecule generation, multiple valid molecules may fulfill the textual requirements (right of the arrow).
To facilitate the evaluation of the above tasks, we also introduce a robust evaluation system with tailored examination process, enabling a comprehensive assessment of LLMs’ performance in open-domain natural language-driven molecule discovery. Furthermore, we propose OpenMolIns, a large-scale instruction-tuning dataset, comprising up to 120 million instruction–molecule pairs, enabling Llama3.1-8B to achieve the state-of-the-art performance among 28 powerful LLMs on S2-Bench.
@@
    We introduce OpenMolIns, an instruction-tuning dataset with up to 120 million instruction–molecule pairs, enabling Llama3.1-8B to achieve SOTA performance on S2-Bench.
@@
    We provide insightful findings based on our extensive benchmarking of 28 LLMs, revealing the limitations of existing targeted generation datasets and highlighting the potential for LLMs to transition from simple pattern recall to realistic molecular design.
2 Speak-to-Structure
@@ 2.1 Task Composition
MolEdit assesses an LLM’s foundational molecular syntax and structural modification capabilities. In this task, a base molecule is provided, and the model is asked to perform a specific change on the molecular structure. In a real-world scenario, this task is analogous to the iterative lead optimization process in drug discovery, where chemists make small, precise changes to a candidate molecule to change its properties. We designed three subtasks to probe this capability: AddComponent and DelComponent test the model’s ability to add or remove a specific functional group, respectively, while SubComponent challenges it to perform a combined operation of both.
MolOpt evaluates an LLM’s ability to perform goal-oriented chemical reasoning. The challenge here is twofold: the LLM should not only edit a molecule but also ensure that the modification leads to a desired change in a specific property. This task directly mirrors the hit-to-lead optimization stage, where molecules are refined to achieve better pharmacological profiles. To assess this, we chose three key properties (LogP, MR, and QED), which are vital for drug discovery and can be reliably and reproducibly calculated. By requiring the model to generate a molecule that improves on a given property, we test its ability to understand the complex relationship between molecular structure and function, a capability that cannot be evaluated by tasks with a single, predetermined answer.
@@ 2.2 Data Construction
Specifically, we design a systematic and scalable data construction process, as illustrated in Figure 4.
For MolEdit and MolOpt, we sample molecules from large, publicly available chemical databases. We selected Zinc-250K (Sterling & Irwin, 2015) for building our test set due to its manageable size and diversity, while the massive PubChem (Kim et al., 2019) database (with over 10 million molecules) serves as the basis for the instruction-tuning dataset. We utilize the RDKit (Landrum, 2013) toolbox to automatically extract key molecular statistics, including structural patterns and properties like LogP, MR, and QED. These extracted features are then integrated into our pre-defined, instruction-based prompt templates, which is demonstrated in Appendix D.
For MolCustom, we move beyond existing databases and programmatically generate instructions to test a model’s ability to create novel molecules. For each of the three subtasks, we randomly generate 5,000 instructions that specify a target number and type of atoms, bonds, or functional groups.
@@ 2.3 Evaluation
MolOpt
Table 1: All used metrics are listed here. ’✓’ means that a metric is calculated on a task.
2.3.1 MolEdit & MolOpt Evaluation
@@
Success Rate: This metric evaluates a model’s ability to fulfill the specific molecule generation requirements. We design automated evaluation processes to verify if the generated molecule meets the specified criteria (e.g., correct functional group modification for MolEdit, or desired property optimization for MolOpt). Detailed Implementations of the automated testing process are shown in Appendix E. This metric examines whether the LLM can follow instructions precisely, which is a fundamental requirement for an LLM to serve as a chemist assistant.
Similarity: In open-domain tasks like MolEdit and MolOpt, a high Success Rate alone is insufficient. We must also ensure that the generated molecule is a reasonable modification of the original, rather than a completely new, unrelated structure that happens to satisfy the criteria. To address this, we measure the Tanimoto Similarity between the generated and original molecules using Morgan Fingerprints (Butina, 1999). A high similarity score indicates that the model has performed a rational, localized edit and has not simply generated a different molecule from scratch. The similarity $\delta(m^{g},m^{o})$ between the generated molecule $m^{g}$ and the original molecule $m^{o}$ is calculated as:
$\delta(m^{g},m^{o})=\frac{\left|fp_{m^{g}}\cap fp_{m^{o}}\right|}{\left|fp_{m^{g}}\cup fp_{m^{o}}\right|},$ (1)
where $fp_{m^{g}}$ and $fp_{m^{o}}$ represents their corresponding Morgan Fingerprints. $\left|fp_{m^{g}}\cap fp_{m^{o}}\right|$ is the size of the intersection between their Morgan Fingerprints, while $\left|fp_{m^{g}}\cup fp_{m^{o}}\right|$ is the union.
Validity: This metric ensures that the generated molecules are chemically valid and follow the rules of molecular syntax. Only the molecules that successfully pass the SMILES parser are considered valid molecules. A higher validity means that the model is more familiar with the molecule syntax.
2.3.2 MolCustom Evaluation
@@
Success Rate: This metric measures how well the generated molecules adhere to the high-level structural constraints provided in the prompt (e.g., number of atoms, bonds, or specific functional groups). As the task requires de novo generation, the success rate directly reflects a model’s ability to translate abstract requirements into a valid, concrete molecular structure. Detailed testing process of MolCustom could also be found in Appendix E.
@@
$n(m^{g})=1-\frac{\sum_{m^{k}\in Zinc}\delta(m^{g},m^{k})}{\left|Zinc\right|},$ (2)
@@
2.3.3 Average Weighted Success Rate
@@
$W\!S\!R_{t}=\begin{cases}n_{t}\times S\!R_{t},&t\in\{MolCustom\}\\ \delta_{t}\times S\!R_{t},&t\in\{MolEdit,MolOpt\}\end{cases},$ (3)
@@
2.4 OpenMolIns: Instruction Tuning Dataset
To effectively train and evaluate LLMs on the open-ended challenges posed by S2-Bench, we introduce OpenMolIns, a specialized instruction-tuning dataset derived from the PubChem database. The core philosophy behind OpenMolIns is to provide models with the kind of flexible, one-to-many training examples they need to learn genuine chemical reasoning, rather than simply memorizing specific input-output pairs. To ensure the integrity of our evaluations, we meticulously designed OpenMolIns to have zero overlap with the Zinc-250K, preventing any data leakage that could compromise the validity of our benchmark results.
We built the instruction-tuning dataset programmatically, creating equal numbers of samples (i.e., 5,000) for all nine subtasks. The RDKit toolbox was essential to this process, allowing us to automatically generate data based on objective molecular properties and rules. This contrasts sharply with previous methods that rely on scarce and costly human annotations. Details of the construction process and the pre-defined prompt templates are shown in Appendix D.
@@ 2.4 OpenMolIns: Instruction Tuning Dataset
Table 2: Statisics of S2-Bench and OpenMolIns.
2.5 Statistics
In this section, we provide a detailed overview of S2-Bench and OpenMolIns, as summarized in Tables 2 and 3. The design and scale of our datasets are not arbitrary; they are meticulously crafted to enable a rigorous and nuanced evaluation of LLMs in open-domain molecule generation, which previous, smaller datasets cannot provide.

### v4

# Speak-to-Structure: Evaluating LLMs in Open-domain Natural Language-Driven Molecule Generation

@@ Abstract.
Recently, Large Language Models (LLMs) have demonstrated great potential in natural language-driven molecule discovery. However, existing datasets and benchmarks for molecule-text alignment are predominantly built on one-to-one mappings, measuring LLMs’ ability to retrieve a single, pre-defined answer, rather than their creative potential to generate diverse, yet equally valid, molecular candidates. To address this critical gap, we propose Speak-to-Structure (S2-Bench), the first benchmark to evaluate LLMs in open-domain natural language-driven molecule generation. S2-Bench is specifically designed for one-to-many relationships, challenging LLMs to exhibit genuine molecular understanding and open-ended generation capabilities. Our benchmark includes three key tasks: molecule editing (MolEdit), molecule optimization (MolOpt), and customized molecule generation (MolCustom), each probing a different aspect of molecule discovery. We also introduce OpenMolIns, a large-scale instruction tuning dataset that enables Llama3.1-8B to surpass the most powerful LLMs like GPT-4o and Claude-3.5 on S2-Bench. Our comprehensive evaluation of 31 LLMs shifts the focus from simple pattern recall to realistic molecular design, paving the way for more capable LLMs in natural language-driven molecule discovery. Our codes and datasets are fully accessible through the Github Repository: https://github.com/phenixace/S2-TOMG-Bench and Huggingface Datasets: https://huggingface.co/datasets/phenixace/S2-TOMG-Bench.
Large Language Models, Benchmark, Text-based Molecule Generation, Molecular Design, Open-ended Generation, Instruction Tuning, Bioinformatics
copyright: acmlicensedjournalyear: 2026copyright: ccconference: Proceedings of the 32nd ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2; August 9–13, 2026; Jeju Island, Republic of Korea.booktitle: Proceedings of the 32nd ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2 (KDD 2026), August 9–13, 2026, Jeju Island, Republic of Koreadoi: 10.1145/3770855.3817473isbn: 979-8-4007-2259-2/2026/08ccs: Applied computing Bioinformaticsccs: Applied computing Molecular sequence analysisccs: Computing methodologies Natural language processing
1. Introduction
@@
Molecule discovery plays a pivotal role in various scientific research fields (Liu et al., 2026, 2025; Zhang et al., 2025; Cao et al., 2022; Li et al., 2026b), from pharmaceuticals (Keiser et al., 2010) to materials science (Higuchi et al., 2023). Traditionally, this is a trial-and-error process (Ekins, 2024) that requires extensive experimentation and data analysis (Mattern and Grosser, 2023), often taking over a decade to bring a new drug candidate to market (Lee et al., 2018).
@@
Recently, Large Language Models (LLMs) have demonstrated great potential in molecule discovery (Edwards et al., 2021, 2022; Chitsaz et al., 2025; Wang et al., 2025; Li et al., 2025c) by leveraging their powerful language understanding and robust reasoning abilities (OpenAI, 2023). Molecules can be represented as linearized representations like SMILES (Weininger, 1988) and SELFIES (Krenn et al., 2022), which allows LLMs to process molecules as textual strings, effectively bridging the gap between molecular structures and natural language. Meanwhile, by aligning molecules with textual descriptions (Edwards et al., 2022; Chen et al., 2025), LLMs can assist chemists in interpreting chemical knowledge, suggesting structural modifications, and predicting compound properties (Li et al., 2025a; Zhang et al., 2024), thereby significantly streamlining the molecule discovery process.
@@
While the integration of LLMs holds immense promise, a significant challenge lies in the limitations of current datasets and benchmarks. For example, existing datasets for aligning molecules and texts like ChEBI-20 (Edwards et al., 2022), PubChem324K (Liu et al., 2023), and M3-20M (Guo et al., 2025b), are constructed based on the one-to-one mapping assumption, where each textual description is linked to a single, predefined target molecule due to the convenience of acquiring ground truth data from existing databases. However, the one-to-one mapping presents a key limitation: it does not fully align with the nature of real-world molecule discovery. In practice, multiple distinct molecules can often share the same desired properties or biological activity (Petrone et al., 2012). For instance, a particular pharmaceutical effect is rarely unique to a single compound, and a material with a specific physical property, like tensile strength, can be realized through various molecular structures. This mismatch between the evaluation paradigm and the reality of molecule discovery casts doubt on whether these datasets and benchmarks genuinely assess the capacity of LLMs for creative molecular design, or if they instead inadvertently encourage models to rely on memorization and pattern-matching.
@@
To bridge this critical gap, we introduce a novel benchmark, Speak-to-Structure (S2-Bench), the first benchmark designed to evaluate the open-domain natural language-driven molecule generation capabilities of LLMs. S2-Bench is well structured into three primary tasks, as shown in Figure 1, each targeting a specific, real-world capability essential for drug and materials science:
@@
Molecule Editing (MolEdit) focuses on testing an LLM’s ability to perform precise, localized structural modifications while preserving main structures. In drug discovery, this is analogous to Lead Optimization (Barcelos et al., 2022), which edits the lead compound to create new variants. This requires a deep understanding of chemical grammar, including valency, stereochemistry, and ring systems, ensuring that edits could create valid molecules.
@@
Figure 1. Task illustration of S2-Bench for open domain natural language-driven molecule generation. In contrast to text-based target molecule generation, multiple valid molecules may fulfill the textual requirements (right of the arrow).
To facilitate the evaluation of the above tasks, we introduce an automated evaluation system tailored to open-domain molecule generation, where exact string matching is neither appropriate nor sufficient. Specifically, we use cheminformatics toolkits to (i) verify validity by parsing the generated outputs; (ii) check constraint satisfaction (e.g., functional-group edits for MolEdit, property direction for MolOpt, and atom/bond/functional-group constraints for MolCustom); and (iii) quantify quality via task-specific metrics, including structural similarity to the source molecule for MolEdit/MolOpt and novelty against a reference database for MolCustom. This design enables a scalable and objective assessment of whether a model follows instructions, performs rational local edits, and generates chemically meaningful candidates.
Furthermore, we propose OpenMolIns, a large-scale instruction-tuning dataset comprising up to 120 million instruction-molecule pairs constructed programmatically from large chemical databases. OpenMolIns mirrors the three tasks in S2-Bench and provides diverse prompt templates to improve robustness to phrasing variations.
@@
    We introduce OpenMolIns, an instruction-tuning dataset with up to 1.2 million instruction–molecule pairs, enabling Llama3.1-8B to achieve SOTA performance on S2-Bench.
@@
    We provide insightful findings based on our extensive benchmarking of 31 LLMs, revealing the limitations of existing targeted generation datasets and highlighting the potential for LLMs to transition from simple pattern recall to realistic molecular design.
2. Related Work
Early work in this direction employed transformer-based models such as MolT5 (Edwards et al., 2022), which leveraged large-scale self-supervised learning to produce high-quality SMILES strings from textual prompts. Building on this foundation, models like KV-PLM (Zeng et al., 2022), MoMu (Su et al., 2022), and BioT5 (Pei et al., 2023) integrated molecular graphs and biochemical texts to jointly enhance molecular understanding and generation.
Parallel to these architecture-driven improvements, LLMs such as MolReGPT (Li et al., 2024) and ICMA (Li et al., 2025a) have demonstrated strong in-context learning capabilities, enabling adaptive molecule generation by retrieving and utilizing relevant examples from the provided context. Besides, MolReFlect (Li et al., 2026a), emphasizes fine-grained alignment between molecular structures and natural language descriptions through a teacher–student paradigm, which effectively captures subtle structural–semantic relationships. Moreover, Mol-R1 (Li et al., 2025b) takes a step further by introducing DeepSeek-R1 like Long-CoT reasoning in molecule generation, enabling robust and interpretable natural language-driven molecule generation.
Meanwhile, Nicolaou and Brown (2013) propose the task of multi-objective molecular optimization and Wu et al. (2024) further extend it to textual requirements by using prompt engineering to guide LLM-based molecule optimization. In addition, Ye et al. (2025) incorporate explicit structural constraints into the optimization process, while InstructMol (Zhuang et al., 2025) further proposes an LLM that can understand and design biomolecules following human instructions.
3. Speak-to-Structure (S2-Bench)
@@ 3.1. Task Composition
MolEdit assesses an LLM’s foundational molecular syntax and structural modification capabilities. In this task, a base molecule is provided, and the model is asked to perform a specific change on the molecular structure. In a real-world scenario, this task is analogous to the iterative lead optimization process in drug discovery, where chemists make small, precise changes to a candidate molecule to change its properties. We designed three subtasks to probe this capability: AddComponent and DelComponent test the model’s ability to add or remove a specific functional group, respectively, while SubComponent challenges a combined operation of both.
MolOpt evaluates an LLM’s ability to perform goal-oriented chemical reasoning. The challenge here is twofold: the LLM should not only edit a molecule but also ensure that the modification leads to a desired change in a specific property. This task directly mirrors the hit-to-lead optimization stage, where molecules are refined to achieve better pharmacological profiles. To assess this, we choose three key properties (LogP, MR, and QED), which are vital for drug discovery and can be reliably and reproducibly calculated. By requiring the model to generate a molecule that improves on a given property, we test its ability to understand the complex relationship between molecular structure and function, a capability that cannot be evaluated by tasks with a single, predetermined answer.
@@ 3.2. Data Construction
Specifically, we design a systematic and scalable data construction process, as shown in Figure 2.
For MolEdit and MolOpt, we sample molecules from large, publicly available chemical databases. We selected Zinc-250K (Sterling and Irwin, 2015) for building our test set due to its manageable size and diversity, while the massive PubChem (Kim et al., 2019) database (with over 10 million molecules) serves as the basis for the instruction-tuning dataset. We utilize the RDKit (Landrum, 2013) toolbox to automatically extract key molecular statistics, including structural patterns and properties like LogP, MR, and QED. These extracted features are then integrated into our pre-defined, instruction-based prompt templates.
For MolCustom, we move beyond existing databases and generate instructions to test a model’s ability to create novel molecules. For each of the three subtasks, we randomly generate 5,000 instructions that specify a target number and type of atoms, bonds, or functional groups.
@@ 3.3. Evaluation
Table 1. All used metrics are listed here. ’✓’ means that a metric is calculated on a task.
@@
3.3.1. MolEdit & MolOpt Evaluation
@@
Success Rate: This metric evaluates a model’s ability to fulfill the specific molecule generation requirements, with values ranging from 0 to 1. We design automated evaluation processes to verify if the generated molecule meets the specified criteria (e.g., correct functional group modification for MolEdit, or desired property optimization for MolOpt). This metric examines whether the LLM can follow instructions precisely, which is a fundamental requirement for an LLM to serve as a chemist assistant.
Similarity: In open-domain tasks like MolEdit and MolOpt, a high Success Rate alone is insufficient. We must also ensure that the generated molecule is a reasonable modification of the original, rather than a completely new, unrelated structure that happens to satisfy the criteria. To address this, we measure the Tanimoto Similarity between the generated and original molecules using Morgan Fingerprints (Butina, 1999). A high similarity score indicates that the model has performed a rational, localized edit and has not simply generated a different molecule from scratch. The similarity $\delta(m^{g},m^{o})\in[0,1]$ between the generated molecule $m^{g}$ and the original molecule $m^{o}$ is calculated as:
(1) $\delta(m^{g},m^{o})=\frac{\left|fp_{m^{g}}\cap fp_{m^{o}}\right|}{\left|fp_{m^{g}}\cup fp_{m^{o}}\right|},$
where $fp_{m^{g}}$ and $fp_{m^{o}}$ represent their corresponding Morgan Fingerprints. $\left|fp_{m^{g}}\cap fp_{m^{o}}\right|$ is the size of the intersection between their Morgan Fingerprints, while $\left|fp_{m^{g}}\cup fp_{m^{o}}\right|$ is the union.
Validity: This metric evaluates the percentage of the generated molecules that are chemically valid and follow the rules of molecular syntax. Molecules need to successfully pass the SMILES parser, and a higher validity means that the model is more familiar with the molecule syntax.
3.3.2. MolCustom Evaluation
@@
Success Rate: This metric measures how well the generated molecules adhere to the high-level structural constraints provided in the prompt (e.g., number of atoms, bonds, or specific functional groups). As the task requires de novo generation, the success rate directly reflects a model’s ability to translate abstract requirements into a valid, concrete molecular structure.
@@
(2) $n(m^{g})=1-\max_{m^{k}\in\text{Zinc}}\delta(m^{g},m^{k})$
@@
3.3.3. Average Weighted Success Rate
@@
(3) $W\!S\!R_{t}=\begin{cases}n_{t}\times S\!R_{t},&t\in\{MolCustom\}\\ \delta_{t}\times S\!R_{t},&t\in\{MolEdit,MolOpt\}\end{cases},$
@@
3.4. OpenMolIns: Instruction Tuning Dataset
Figure 2. Data construction workflow of S2-Bench & OpenMolIns.
To effectively train and evaluate LLMs on the open-ended challenges posed by S2-Bench, we introduce OpenMolIns, a specialized instruction-tuning dataset derived from the PubChem database. The core philosophy behind OpenMolIns is to provide models with the kind of flexible, one-to-many training examples they need to learn genuine chemical reasoning, rather than simply memorizing specific input-output pairs. To ensure the integrity of our evaluations, we meticulously designed OpenMolIns to have zero overlap with Zinc-250K, preventing any data leakage that could compromise the validity of our benchmark results. Detailed Comparison of the data construction process between S2-Bench and OpenMolIns is shown in Figure 2.
We built the instruction-tuning dataset programmatically, creating equal numbers of samples (i.e., 5,000) for all nine subtasks. The RDKit toolbox was essential to this process, allowing us to automatically generate data based on objective molecular properties and rules. This contrasts sharply with previous methods that rely on scarce and costly human annotations.
@@ 3.4. OpenMolIns: Instruction Tuning Dataset
3.5. Statistics
Here, we provide a detailed overview of S2-Bench and OpenMolIns, as summarized in Tables 2 and 3. The design and scale of our datasets are not arbitrary; they are meticulously crafted to enable a rigorous and nuanced evaluation of LLMs in open-domain molecule generation, which previous, smaller datasets cannot provide.